# Search Arena: Analyzing Search-Augmented LLMs

### v1

@@ 3.3 Cross-Setting Analysis
Search Arena evaluates models in a setting where user prompts and expectations are conditioned on models’ access to web search. In this section, we study how search and non-search models perform under different prompt distributions and user expectations–specifically Search Arena vs Text Arena. Additional results on model performance changes across different benchmarks are provided in Appendix F.
To investigate performance differences across settings, we deployed Gemini-2.5 Pro Experimental [17]–with and without access to web search–to the Search and Text Arenas [10]. In the Text Arena, the search model only competed against its non-search version; in the Search Arena, the non-search model competed against all other supported search models. Additionally, inline citations were disabled for the search model in the Text Arena to avoid vote bias. In the Text Arena, users typically assume that models operate in closed-book settings without access to external information, while in the Search Arena, user expectations are explicitly conditioned on the search setting. This setup enables us to examine whether and under what conditions access to web search enhances or degrades model performance.
@@
[…]
These cross-arena experiments demonstrate that search-augmentation does not hurt performance in non-search settings and can improve responses to queries related to information retrieval and synthesis. However, removing web search significantly hurts model performance in search settings.
@@ 4 Related Work
Large Language Models. LLMs have made impressive advances in language understanding, dialogue generation, and reasoning, enabled by techniques such as large-scale pretraining [2, 3, 6, 18, 36, 56], chain-of-thought prompting [31, 59, 51, 58, 66], and reinforcement learning with human feedback (RLHF) [4, 43]. The dataset and evaluation landscape has progressed from standardized, static evaluations [22, 28, 23, 70] toward more challenging settings such as deep reasoning [35, 30, 68], coding [27, 26], and open-ended dialogue under crowd-sourced evaluation [10, 70]. While crowd-sourced setups can mitigate concerns over data contamination in pretraining [64, 9], recent work has also highlighted potential biases and oversights in human preferences [12, 62].
@@
As LLMs gain tool-use capabilities (e.g., APIs, code interpreters, and web browsers) [24, 49, 18, 46, 67], domain-specific benchmarks and datasets have emerged to analyze and evaluate LLMs under different environments [73, 34, 38]. In the context of web search, several search-augmented LLMs have been developed [41, 47, 16], which retrieve live information to support better reasoning. However, existing benchmarks such as SimpleQA [60] and BrowseComp [61] are limited to single-turn, fact-based, monolingual queries. Although WebArena [73] contains diverse user prompts and a web-based interface, it emphasizes closed-world web navigation tasks rather than open-ended search, reasoning, and dialogue. We introduce Search-Arena, the first large-scale, crowd-sourced dataset for search-augmented LLMs with human preference signals, covering diverse intents, topics, and multi-turn interactions across 70+ languages, collected through a transparent, open platform.
@@
Traditional and LLM-Integrated Information Retrieval (IR). Information retrieval is a long-standing task, with early methods like BM25 [48], PageRank [45], and embedding-based approaches [39, 25]. In the IR field, several static benchmarks [54, 40] and large-scale web search datasets with user logs and preference signals [57, 13, 8] have enabled robust evaluation of retrieval systems and comprehensive studies on user behavior.
@@
With the rise of LLMs, information retrieval has moved beyond traditional search and become integrated into LLM workflows. In Retrieval Augmented Generation (RAG) settings, retrieved text is appended to the input prompt [32, 20, 24, 11]; later work explored search-augmented systems operating in the open web [24, 49, 67]. Datasets for evaluating these LLM-IR systems span needle-in-a-haystack retrieval [29, 63], citation attribution [69, 1], and general question answering [65, 21, 60, 61]. For search augmented LLMs, existing benchmarks lack full human–AI interaction traces, including queries, retrieved documents, responses, and preferences, which are needed for human-centric analysis as done in traditional IR [13, 8]. To fill this gap, we introduce Search Arena, a large-scale open dataset that enables human-centered analysis of this emerging interface.
@@ 5 Limitations, Conclusion, and Future Work
Limitations and Broader Impact. Crowd-sourced data analysis provides valuable insight into real-world user preferences but comes with limitations and broader social implications. First, the collected data may reflect demographic skews and may not be fully representative of the broader population, as not all users choose to vote and human judgments are inherently subjective [7, 12]. Second, because conversational data with human preferences is personal and potentially valuable for model improvement [44], its release requires careful consideration of both privacy and equitable access across the community.
To address these concerns, we anonymized model responses and randomized their left/right placement to reduce known human biases. During the data collection period, no early access to data was granted to model providers, nor were any pre-release models deployed on the platform. We obtained user consent at interaction time and enforced a strict privacy policy. To aid responsible interpretation, we also analyzed known biases (e.g., response length; Subsection 3.1) and reported user demographics in Figure A4 and Figure A8. Additional details are provided in Appendix A.
Conclusions and Future Work. We present Search Arena, the first large-scale dataset and analysis of human interactions with search-augmented LLMs. Spanning over 24k multi-turn conversations and 12k human preference votes across more than 70 languages, the dataset captures a broader range of user intents and topics than prior benchmarks. Our analysis reveals that user preference is positively associated with citation count and certain source types; however, models do not always cite correctly, highlighting a key challenge in trustworthy systems. The cross-arena experiment further shows that search-augmented and non-search models behave differently across settings. We release the full dataset to support future research in search-augmented LLMs and human-centric analyses.
@@ Acknowledgments
  • Appendix A describes the platform design, model configurations, and the limitations and policies of our data collection process.
  • Appendix B outlines the data processing steps, including intent classification, topic modeling, and filtering procedures.
  • Appendix C includes details (e.g., parameters, prompts, raw outputs) on the dataset differencing experiments.
@@
  • Appendix E reports citation analysis, including user preference experiments and implementation details of the citation attribution pipeline.
@@
Figure A1: Search Arena User Interface. Search Arena is integrated into the Chatbot Arena ecosystem, sharing the same front-end design and entry point. As shown at the bottom, it is implemented as a separate tab. The central panel displays a side-by-side chat interface with two anonymous models, each providing responses that include clickable inline citations and expandable reference links. Users can engage in multi-turn conversations and cast a vote at any time during the conversation using the four feedback options (A is better, B is better, Tie, and Both are bad).
@@ Appendix A Search Arena Platform
As described in Section 2, Search Arena (https://legacy.lmarena.ai/?search) is an open, crowdsourced evaluation platform for search-augmented LLMs, launched on March 18, 2025. This section outlines our data collection and release protocols, followed by a description of the supported models and key design decisions.
@@ A.1 Data
The Search Arena platform does not require user login, but users must explicitly accept the Terms of Service before using the platform. Once accepted, users access a dedicated tab within Chatbot Arena to interact with search-enabled models. This setup naturally leads to more information-seeking queries compared to the default Text Arena interface. For each prompt, responses from two anonymous models are displayed side-by-side, and users may cast a preference vote at any time during the interaction. The user interface is shown in Figure A1. The full text of Terms of Service is at Figure A2.
Over the course of a 7-week data collection period, the platform recorded an average of roughly 800-1,500 conversations per day. After filtering out examples with server errors, inconsistent configurations, or other quality issues, we retain approximately 24,000 conversations, about half of which include human preference votes. Figure A3 shows daily traffic trends, with spikes aligned to major model updates or platform announcements, and dips on weekends.
@@
Figure A2: Term of Service of the Chatbot Arena Platform (including Search Arena).
Figure A3: Daily traffic on the Search Arena platform.
@@
Per our Terms of Service (Figure A2), all conversation data is released under a Creative Commons Attribution (CC-BY) license. To protect user privacy, we apply automated de-identification using Google’s Data Loss Prevention (DLP) API, which removes personal identifiers such as email addresses, credit card numbers, and API keys. Approximately 2% of examples were flagged by this process.
@@ A.2 Models
Search Arena currently supports 12 search-augmented LLMs from Perplexity, Gemini, and OpenAI, as summarized in Table A1. Unless otherwise noted, we group models with different inline citation styles (elaborated below) of the same base model. Each model is accessed via its provider’s public API using the default configuration, including search_context_size=medium for Perplexity and OpenAI models. We also evaluate several additional variants, including (1) OpenAI’s gpt-4o-search-preview with user_location enabled and (2) Perplexity and OpenAI models configured with search_context_size=high. Figure A9 shows the number of battles per model.
@@
Table A1: List of supported models and their configurations on Search Arena. We evaluate OpenAI’s web search API, which differs from the search feature in the ChatGPT product.
@@ B.1 Linguistic Features
Users’ demographic information in Search Arena, based on country codes extracted from IP addresses, is shown in Figure A4. The dataset includes 11,650 unique users across 136 countries, with the United States (18.7%), Russia (9.9%), Germany (5.9%), and China (5.5%) as the top four.
Language distribution (of the top 50 languages) of Search Arena prompts is shown in Figure A8. The prompts span 71 languages with English (56.4%), Russian (11.4%), and Chinese (6.8%) as the top three.
As shown in Figure A5, Search Arena prompt lengths vary across intent categories. As expected, Text Processing (85.1 words) and Analysis (72.7 words) prompts are generally longer, while Explanation (24.7 words) and Factual Lookup (16.3 words) prompts are shorter.
Search Arena conversations are multi-turn, with 22.4% of all conversations containing more than one turn. Specifically, there are 3,288 conversations with 2 turns, 966 with 3 turns, and 460 with 4 turns. The distribution is shown in Figure A6.
@@ B.2 User Intent Annotation
We then scaled the annotation to the full dataset using GPT-4.1 [42], using the same in-context examples from LABEL:tab:intent_rubric. The total annotation cost was approximately \$20 USD. To evaluate label quality, we compared GPT-4.1 annotations with human labels on 150 samples drawn from the top three languages in our dataset (English, Russian, and Chinese). The resulting Cohen’s Kappa score of 0.812 on top-2 intents indicates strong agreement. The full GPT-4.1 prompt is provided below:
[…]
@@
Table A2: User intent taxonomy including category descriptions and examples.
@@
Text Processing Users request linguistic tasks such as summarizing, translating, paraphrasing, or refining text. Summarize <website>. Translate this paragraph from Spanish to English. Rephrase this email professionally for an investment-banking application.
@@
Explanation Users seek detailed clarifications, educational insights, or thorough elaborations aimed at better understanding a concept, process, or phenomeneon. May ask “why” or “how” something works without implying action. What led to the fall of the Roman Empire? Can you explain the Heisenberg uncertainty principle? How does gradient-descent optimization work in machine learning?
@@
Figure A4: Search Arena Users’ Demographics. Search Arena data includes 11,650 unique users across 136 countries.
Figure A5: Search Arena Prompt Length Distribution by Intent. Prompt lengths vary across intent categories. Text Processing (85.1 words) and Analysis (72.7 words) prompts are generally longer, while Explanation (24.7 words) and Factual Lookup (16.3 words) prompts are shorter.

### v2

@@ 3.3 Cross-Setting Analysis
Search Arena evaluates models where user prompts and expectations are conditioned on models’ access to web search. In this section, we study how search and non-search models perform under various prompt distributions and user expectations–specifically, Search Arena vs Text Arena. Additional results on model performance across benchmarks are provided in Appendix F.
To investigate performance differences across settings, we deployed Gemini-2.5 Pro Experimental (Gemini, 2025)–with and without access to web search–to the Search and Text Arenas. In the Text Arena, the search model only competed against its non-search version; in the Search Arena, the non-search model competed against all other supported search models. Additionally, inline citations were disabled for the search model in the Text Arena to avoid vote bias. In the Text Arena, users typically assume that models operate in closed-book settings without access to external information, while in the Search Arena, user expectations are explicitly conditioned on the search setting. This setup enables us to examine whether and under what conditions access to web search enhances or degrades model performance.
@@
Text Arena Setting. We collected 544 battles between the search and non-search versions of the model, yielding 245 ties (45%), 143 search-preferred (26%), and 156 non-search preferred votes (28%). We observe a high proportion of tie votes, and the difference between search-preferred and non-search-preferred votes is not statistically significant ($\text{p-value}=0.244$). On aggregate, search and non-search models have comparable performance in the Text Arena. To further analyze performance differences by prompt type, we apply the intent classification pipeline from Subsection 2.2 on the 544 collected Text Arena prompts. The distribution of votes by intent class is shown in Figure 8 (Left). We observe a high proportion of ties in Analysis, Creative Generation, and Guidance queries, indicating that there are no significant differences between model responses for these types of prompts. However, for Factual Lookup ($\text{p-value}=0.012$) and Info Synthesis ($\text{p-value}=0.095$), the difference is more pronounced in favor of the search model. Thus, the search-augmented model is preferred for knowledge acquisition tasks–even in the absence of well-defined user expectations–because web-grounded responses typically provide precise data, statistics, dates, names, and domain-specific terminology. We also note that the difference in performance for Text Processing ($\text{p-value}=0.077$) queries favors the non-search model. In these cases, the non-search model often provides structured responses (e.g., numbered or bulleted lists, headings), which suggests that presentation style may impact user evaluations.
Search Arena Setting. We collected 315 pairwise battles between a non-search model and search-augmented models in Search Arena, with 99 ties (31%), 126 search-preferred (40%), and 90 non-search-preferred votes (29%). The difference between search and non-search-preferred votes is statistically significant ($\text{p-value}=0.009$); the non-search model underperforms under a search-conditioned distribution. A detailed vote distribution by user intent is shown in Figure 8 (Right). Compared with the Text Arena (Figure 8 (Left)), tie votes are less frequent across all categories, indicating that the differences between model responses are more pronounced. The difference is most expressed for Factual Lookup ($\text{p-value}=5.8\times 10^{-5}$) and Info Synthesis ($\text{p-value}=0.092$) queries.
These cross-arena experiments demonstrate that search augmentation does not hurt performance in non-search settings and can improve responses to queries related to information retrieval and synthesis. However, removing web search significantly hurts model performance in search settings.
@@ 4 Related Work
Large Language Models. LLMs have made impressive advances in language understanding, dialogue generation, and reasoning, enabled by techniques such as large-scale pretraining (Anthropic, 2024; Bai et al., 2023; Brown et al., 2020; Grattafiori et al., 2024; Liu et al., 2024a; Touvron et al., 2023), chain-of-thought prompting (Kojima et al., 2023; Wei et al., 2023; Shinn et al., 2023; Wang et al., 2023; Yao et al., 2023a), and reinforcement learning with human feedback (RLHF) (Bai et al., 2022; Ouyang et al., 2022a). The dataset and evaluation landscape has moved from static benchmarks (Hendrycks et al., 2021a; Joshi et al., 2017; Hendrycks et al., 2021b) toward more challenging settings such as deep reasoning (Li et al., 2024; Kazemi et al., 2025; Zeng et al., 2024; Lin et al., 2025), coding (Jimenez et al., 2024; Jain et al., 2024), and open-ended dialogue under crowd-sourced evaluation (Chiang et al., 2024; Zhao et al., 2024). While crowd-sourced setups can mitigate concerns over data contamination in pretraining (Yang et al., 2023; Cheng et al., 2025b), recent work shows potential biases and oversights in human preferences (Clark et al., 2021; Wu & Aji, 2025).
@@
As LLMs gain tool-use capabilities (e.g., APIs, code interpreters, and web browsers) (Hilton et al., 2021; Schick et al., 2023; Grattafiori et al., 2024; Patil et al., 2024; Yao et al., 2023b), domain-specific benchmarks and datasets have emerged to analyze and evaluate LLMs under different environments (Zhou et al., 2024; Li et al., 2023b; Liu et al., 2023). In the context of web search, several search-augmented LLMs have been developed (OpenAI, 2024; Perplexity, ; Gemini, 2024), which retrieve live information to support better reasoning. However, existing benchmarks such as SimpleQA (Wei et al., 2024) and BrowseComp (Wei et al., 2025) are limited to single-turn, fact-based, monolingual queries. Although WebArena (Zhou et al., 2024) contains diverse user prompts and a web-based interface, it emphasizes closed-world web navigation tasks rather than open-ended search, reasoning, and dialogue. We introduce Search-Arena, the first large-scale, crowd-sourced dataset for search-augmented LLMs with human preference signals, covering diverse intents, topics, and multi-turn interactions across 70+ languages, collected through a transparent, open platform.
@@
Traditional and LLM-Integrated Information Retrieval (IR). Information retrieval is a long-standing task, with early methods like BM25 (Robertson et al., 1994), PageRank (Page et al., 1998), and embedding-based approaches (Monir et al., 2024; Huang et al., 2020). Several static benchmarks (Thakur et al., 2021; Muennighoff et al., 2023) and large-scale web search datasets with user logs and preference signals (Voorhees & Harman, 2005; Craswell et al., 2020; Chen et al., 2024b) have enabled robust evaluation of retrieval systems and comprehensive studies on user behavior.
@@
With the rise of LLMs, information retrieval has moved beyond traditional search and become integrated into LLM workflows. In Retrieval Augmented Generation (RAG) settings, retrieved text is appended to the input prompt (Lewis et al., 2021; Guu et al., 2020; Hilton et al., 2021; Chuang et al., 2025); Datasets for evaluating these LLM-IR systems span needle-in-a-haystack retrieval (Kamradt, 2023; Wu et al., 2025), citation attribution (Zhang et al., 2024; Abolghasemi et al., 2024), and general question answering (Yang et al., 2018; Han et al., 2024; Wei et al., 2024; 2025). More recently, this line of work has been extended from direct single-round information-seeking to natural, conversational search scenarios (Gao et al., 2023; Mo et al., 2025; Cheng et al., 2025a; Adlakha et al., 2022), and then to search-augmented LLM systems operating over the open web (Hilton et al., 2021; Schick et al., 2023; Yao et al., 2023b). We view search-augmented LLM systems as a variant of RAG, with clear differences, including (1) “unbounded” search space over the whole open web and (2) multi-round agent-based query refinement, search, and generation. However, for search augmented LLMs, existing benchmarks lack full human-AI interaction traces, including queries, retrieved documents, responses, and preferences, which are needed for human-centric analysis as done in traditional IR (Craswell et al., 2020; Chen et al., 2024b). To fill this gap, we introduce Search Arena, a large-scale open dataset that enables human-centered analysis of this emerging interface.
@@ 5 Limitations, Conclusion, and Future Work
Limitations, Future Work, and Broader Impact. Crowd-sourced data analysis provides valuable insight into real-world user preferences but comes with limitations and broader social implications. First, the collected data may reflect demographic skews and may not be fully representative of the broader population, as not all users choose to vote and human judgments are inherently subjective (Chen et al., 2024a; Clark et al., 2021). Second, because conversational data with human preferences is personal and potentially valuable for model improvement (Ouyang et al., 2022b), its release requires careful consideration of both privacy and equitable access across the community.
To address these concerns, we anonymized model responses and randomized their left/right placement to reduce known human biases. During the data collection period, no early access to data was granted to model providers, nor were any pre-release models deployed on the platform. We obtained user consent at interaction time and enforced a strict privacy policy. To aid responsible interpretation, we also analyzed known biases (e.g., response length; Subsection 3.1) and reported user demographics in Figure B.1 and Figure B.5. Additional details are provided in Appendix A.
We identify several limitations in our current study that open avenues for future research. While our primary focus is dataset construction and interaction analysis, the Search Arena dataset facilitates the following directions:
  • Objective metrics: As detailed in Section 3, user preferences are often influenced by surface-level features, such as citation counts or source reputation, rather than the intrinsic quality of the cited sources and the generated content. Consequently, user-provided labels may diverge from objective standards like factuality, citation-claim consistency, and source specificity. Future work should expand the evaluation framework to include these objective metrics and analyze their correlation with human preference signals to better understand the gap between perceived credibility and actual groundedness.
  • Offline evaluation: Our current evaluation relies on an online human-in-the-loop setup, which, while robust, restricts the scalability of testing new models. Future work can leverage our dataset to construct automated benchmarks for offline evaluation. This could involve training reward models or LLMs that approximate human preferences, potentially combining them with the objective metrics discussed above to create a comprehensive scoring rubric.
  • Model development: Our cross-arena and preference analyses reveal that blindly augmenting models with search is not always optimal. Future research can utilize these insights to improve both the search module (e.g., better source filtering) and the backbone LLM (e.g., optimizing policies for when to trigger search versus relying on parametric knowledge).
Conclusion. We present Search Arena, the first large-scale dataset and analysis of human interactions with search-augmented LLMs. Spanning over 24k multi-turn conversations and 12k human preference votes across more than 70 languages, the dataset captures a broader range of user intents and topics than prior benchmarks. Our analysis reveals that user preference is positively associated with citation count and certain source types; however, models do not always cite correctly, highlighting a key challenge in trustworthy systems. The cross-arena experiment further shows that search-augmented and non-search models behave differently across settings. We release the full dataset to support future research in search-augmented LLMs and human-centric analyses.
@@ Acknowledgments
Appendix A describes platform design, supported model, and policies of our data collection process.
Appendix B outlines intent classification, topic modeling, and filtering procedures.
Appendix C includes details (e.g., parameters, prompts) on the dataset differencing experiments.
@@
Appendix E outlines citation analysis, user preference experiments, and citation attribution pipeline.
@@
Appendix G includes comparative analysis with additional related datasets.
@@ Appendix A Search Arena Platform
Figure A.1: Search Arena User Interface. Search Arena is integrated into the Chatbot Arena ecosystem, sharing the same front-end design and entry point. As shown at the bottom, it is implemented as a separate tab. The central panel displays a side-by-side chat interface with two anonymous models, each providing responses that include clickable inline citations and expandable reference links. Users can engage in multi-turn conversations and cast a vote at any time during the conversation using the four feedback options (A is better, B is better, Tie, and Both are bad).
As described in Section 2, Search Arena (https://arena.ai/search) is an open, crowdsourced evaluation platform for search-augmented LLMs, launched on March 18, 2025. It should be noted that the platform’s user interface has been changed and the data for this study has been collected through the legacy interface (see Figure A.1). This section outlines our data collection and release protocols, followed by a description of the supported models and key design decisions.
@@ A.1 Data
The Search Arena platform does not require user login, but users must explicitly accept the Terms of Service before using the platform. Once accepted, users access a dedicated tab within Chatbot Arena to interact with search-enabled models. This setup naturally leads to more information-seeking queries compared to the default Text Arena interface. For each prompt, responses from two anonymous models are displayed side-by-side, and users may cast a preference vote at any time during the interaction. The user interface is shown in Figure A.1. The full text of Terms of Service is at Figure A.2.
Over the course of a 7-week data collection period, the platform recorded an average of roughly 800-1,500 conversations per day. After filtering out examples with server errors, inconsistent configurations, or other quality issues, we retain approximately 24,000 conversations, about half of which include human preference votes. Figure A.3 shows daily traffic trends, with spikes aligned to major model updates or platform announcements, and dips on weekends.
@@
Figure A.2: Term of Service of the Search Arena Platform.
Figure A.3: Daily traffic on the Search Arena platform.
@@
Per our Terms of Service (Figure A.2), all conversation data is released under a Creative Commons Attribution (CC-BY) license. To protect user privacy, we apply automated de-identification using Google’s Data Loss Prevention (DLP) API, which removes personal identifiers such as email addresses, credit card numbers, and API keys. Approximately 2% of examples were flagged by this process. Users may also contact the platform provider to request their records be removed from the released dataset through privacy@arena.ai. Details on the privacy policy of the platform are provided here.
@@ A.2 Models
Search Arena currently supports 12 search-augmented LLMs from Perplexity, Gemini, and OpenAI, as summarized in Table A.1. Unless otherwise noted, we group models with different inline citation styles (elaborated below) of the same base model. Each model is accessed via its provider’s public API using the default configuration, including search_context_size=medium for Perplexity and OpenAI models. We also evaluate several additional variants, including (1) OpenAI’s gpt-4o-search-preview with user_location enabled and (2) Perplexity and OpenAI models configured with search_context_size=high. Figure D.1 shows the number of battles per model.
@@
Table A.1: List of supported models and their configurations on Search Arena. We evaluate OpenAI’s web search API, which differs from the search feature in the ChatGPT product.
@@ B.1 Linguistic Features
Users’ demographic information in Search Arena, based on country codes extracted from IP addresses, is shown in Figure B.1. The dataset includes 11,650 unique users across 136 countries, with the United States (18.7%), Russia (9.9%), Germany (5.9%), and China (5.5%) as the top four.
Language distribution (of the top 50 languages) of Search Arena prompts is shown in Figure B.5. The prompts span 71 languages with English (56.4%), Russian (11.4%), and Chinese (6.8%) as the top three.
As shown in Figure B.2, Search Arena prompt lengths vary across intent categories. As expected, Text Processing (85.1 words) and Analysis (72.7 words) prompts are generally longer, while Explanation (24.7 words) and Factual Lookup (16.3 words) prompts are shorter.
Search Arena conversations are multi-turn, with 22.4% of all conversations containing more than one turn. Specifically, there are 3,288 conversations with 2 turns, 966 with 3 turns, and 460 with 4 turns. The distribution is shown in Figure B.3.
@@ B.2 User Intent Annotation
We then scaled the annotation to the full dataset using GPT-4.1 (OpenAI, 2025), using the same in-context examples from LABEL:tab:intent_rubric. The total annotation cost was approximately \$20 USD. To evaluate label quality, we compared GPT-4.1 annotations with human labels on 150 samples drawn from the top three languages in our dataset (English, Russian, and Chinese). The resulting Cohen’s Kappa score of 0.812 on top-2 intents indicates strong agreement. The full GPT-4.1 prompt is provided below.
Prompts for LLM-based Intent Classification You are an impartial classifier.
@@
Table B.1: User intent taxonomy including category descriptions and examples.
@@
Text Processing Users request linguistic tasks such as summarizing, translating, paraphrasing, or refining text. Summarize <website>. Translate this paragraph from Spanish to English. Rephrase this email professionally for an investment‑banking application.
@@
Explanation Users seek detailed clarifications, educational insights, or thorough elaborations aimed at better understanding a concept, process, or phenomeneon. May ask “why” or “how” something works without implying action. What led to the fall of the Roman Empire? Can you explain the Heisenberg uncertainty principle? How does gradient‑descent optimization work in machine learning?
@@
Figure B.1: Search Arena Users’ Demographics. Search Arena data includes 11,650 unique users across 136 countries.
Figure B.2: Search Arena Prompt Length Distribution by Intent. Prompt lengths vary across intent categories. Text Processing (85.1 words) and Analysis (72.7 words) prompts are generally longer, while Explanation (24.7 words) and Factual Lookup (16.3 words) prompts are shorter.